OpenStudy (anonymous):

A man pulls a box with a rope parallel to the ground. He pulls the box 100.0 m and does 5,000.0 J of work. What was the mass of the box? 0.5 kg 5 kg-my answer 50 kg 500 kg

OpenStudy (amistre64):

Work = force times distance, units of N.m 5000 = F * 100 50N = F N = mass * m^2/s^2 .... can we assume thats mg? 50/abt10 = 5 sounds right to me ..... but its a little hazy

OpenStudy (anonymous):

Work = Force * Distance \[Force = \frac{ 5000 Joules }{ 100m } = 50 Newton\] Force = Mass * Gravity \[Mass = \frac{ Force }{ Gravitational-acceleration } = \frac{ 50 Newtons }{ 9.81\frac{ m }{ s^2 } } = 5.0968 kg\] So 5 kg would be correct.